OpenStudy (anonymous):

y''+3y'-10y=t+e^(-t) . . . I can find the complimentary function but am struggling to find the particular integral! Any help appreciated!:-D

OpenStudy (amistre64):

whats the comp?

OpenStudy (amistre64):

it also might be helpful to know what methods you have been taught there is a table method (trial and error really) there is the variation of parameters and a wronskin

OpenStudy (anonymous):

Ae^(-4x) + Be^(2x)

OpenStudy (amistre64):

the long version, which i learned on ... is to assume, and since your setup is wrt "t" and not "x" \[y_h=Ae^{-4t} + Be^{2t}\]and take the derivatives, assuming A and B are functions of t

OpenStudy (amistre64):

what is y'h?

OpenStudy (amistre64):

make that yp :) im used to homogenous (yh) for "compliment"

OpenStudy (anonymous):

Thats the part I'm struggling with. Yp=Ct+D I think!

OpenStudy (amistre64):

yp, not yh, to fix any ambiguity in this \[y_p=Ae^{-4t} + Be^{2t}\] \[y'_p=-4Ae^{-4t}+A'e^{-4t}+ 2Be^{2t} + B'e^{2t}\] let A'+B' = 0 to account for the homogenous results \[y'_p=-4Ae^{-4t}+2Be^{2t}\] \[y''_p=16Ae^{-4t}-4'Ae^{-4t}+4Be^{2t}+2B'e^{2t}\] now insert these original setup into the

OpenStudy (amistre64):

...insert those into the original setup :/

OpenStudy (amistre64):

y''+3y'-10y \[~~~~~~~y''_p:~~16Ae^{-4t}-4'Ae^{-4t}+4Be^{2t}+2B'e^{2t}\\ ~~~~~~3y'_p:-12Ae^{-4t}~~~~~~~~~~~~~~~+6Be^{2t}\\ -10y_p:-10Ae^{-4t} ~~~~~~~~~~~~~~-10 Be^{2t}\] well the Bs are nice enough to zero out for us :/ whats up with those As?

OpenStudy (amistre64):

am i going to have to dbl chk your characteristic?

OpenStudy (amistre64):

y''+3y'-10y r^2 + 3r - 10 = 0 r = -2 and 5

OpenStudy (amistre64):

sigh .... why would youwant to lie to me?

OpenStudy (anonymous):

apologies, got mixed up between to different examples!

OpenStudy (amistre64):

\[~~~~~~~y''_p:~~~25Ae^{5t}+5A'e^{5t}+4Be^{2t}-2B'e^{-2t}\\ ~~~~~~3y'_p:~~~15Ae^{5t}~~~~~~~~~~~~~~~-6Be^{-2t}\\ -10y_p:-10Ae^{5t} ~~~~~~~~~~~~~~-10 Be^{-2t}\] is that fixed better?

OpenStudy (amistre64):

got me signs mixed, roots are -5 and 2 :)

OpenStudy (amistre64):

\[~~~~~~~y''_p:~~~25Ae^{-5t}-5A'e^{5t}+4Be^{2t}+2B'e^{2t}\\ ~~~~~~3y'_p:-15Ae^{-5t}~~~~~~~~~~~~~~~+6Be^{2t}\\ -10y_p:-10Ae^{-5t} ~~~~~~~~~~~~~~-10 Be^{2t}\\----------------------\\ t+e^{-t}=~~~-5A'e^{5t}+2B'e^{2t}\\\]\[0=A'e^{-5t}+B'e^{2t}\] now we have a system of equations

OpenStudy (amistre64):

this system presents a Cramer rule that is equivalent to the Wronskian method